Circular No. 8964

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2008 O3 (BOATTINI) A. Boattini, University of Arizona, reports his discovery of a comet

on CCD images taken with the Mt. Lemmon 1.5-m reflector (discovery observation tabulated below); four co-added 24-s exposures show moderate condensation, a coma diameter of 5′′–6′′, and a faint tail ∼ 30′′ long in p.a. 250^{o}–255^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, S. Foglia and L. Buzzi report that CCD images taken by R. Holmes with a 0.61-m telescope at Charleston, IL, U.S.A., on Aug. 1.4 UT show a diffuse coma of diameter 7′′, elongated in p.a. 235^{o}.

2008 UT  α_{2000}  δ_{2000}  Mag.
July 29.46129  4^{h}20^{m}33.^{s}03  +21^{o}01′34′′.8  19.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2008-P05.

T = 2008 June 1.546 TT  ω = 339.^{o}806
Ω = 48.343 }2000.0
q = 2.56893 AU  i = 31.342

COMETS C/2008 J7–J12 (SOHO) Further to IAUC 8962, additional apparent comets have been found

on SOHO website images, all Kreutz sungrazers except for C/2008 J10 and C/2008 J12 (Meyer group); finder code GP = G. Pappa. C/2008 J7 was very small and slightly diffuse (mag 7.5–8). C/2008 J8 was very diffuse and very faint (mag ∼ 8). C/2008 J9 was small and condensed (mag 7–7.5). C/2008 J10 was small and essentially stellar in appearance (mag ∼ 7.5). C/2008 J11 was very small and slightly diffuse (mag ∼ 8). C/2008 J12 was stellar in appearance; in C2 images it was of mag ∼ 6.5, while in C3 it appeared very small and faint.

Comet 2008 UT α_{2000} δ_{2000} Inst. F MPEC
C/2008 J7 May 1.631 2^{h}39.^{m}7 +13^{o}50′ C2  HS 2008-L48
C/2008 J8  5.038 2 53.8 +14 54 C2  JR 2008-L48
C/2008 J9  5.338 2 55.6 +15 01 C2  MK 2008-L48
C/2008 J10  6.979 3 01.5 +17 34 C2  JR 2008-M10
C/2008 J11  7.588 3 04.3 +15 39 C2  GP 2008-M10
C/2008 J12  7.963 3 05.3 +17 32 C2/3 SY 2008-M10
2008 August 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2008 CBAT  Daniel W. E. Green